# Adapt On-the-Go: Behavior Modulation for Single-Life Robot Deployment

### v2

@@ Abstract
To succeed in the real world, robots must cope with situations that differ from those seen during training. We study the problem of adapting on-the-fly to such novel scenarios during deployment, by drawing upon a diverse repertoire of previously-learned behaviors. Our approach, RObust Autonomous Modulation (ROAM), introduces a mechanism based on the perceived value of pre-trained behaviors to select and adapt pre-trained behaviors to the situation at hand. Crucially, this adaptation process all happens within a single episode at test time, without any human supervision. We demonstrate that ROAM enables a robot to adapt rapidly to changes in dynamics both in simulation and on a real Go1 quadruped, even successfully moving forward with roller skates on its feet. Our approach adapts over 2x as efficiently compared to existing methods when facing a variety of out-of-distribution situations during deployment by effectively choosing and adapting relevant behaviors on-the-fly.
Keywords: reinforcement learning, single-life deployment, autonomous reinforcement learning, out-of-distribution adaptation
Figure 1: On-The-Go Adaptation via Robust Autonomous Modulation (ROAM). An agent will inevitably encounter a wide variety of situations during deployment, and handling such situations may require a variety of different behaviors. We propose Robust Autonomous Modulation (ROAM), which dynamically employs and adapts relevant behaviors as different situations arise during a single trial of deployment.
@@ 1 Introduction
Existing adaptation methods often operate within the standard, episodic training paradigm where the agent is assumed to be reset and have another chance to attempt the task each time [cully2015robots, song2020rapidly, julian2020never]. However, these are idealized conditions that in practice often rely on human intervention. In the above search-and-rescue example, the robot’s state cannot be arbitrarily restored; during deployment, the robot cannot be repaired, and it is not feasible for a human to fetch it if it gets stuck in a situation it is not equipped to handle. This situation necessitates adapting at test time both quickly and autonomously, to succeed at the task within a single episode. Therefore, we frame our problem setting as an instantiation of single-life deployment [chen2022you], where the agent is given prior knowledge from ‘past lives’ but evaluated on its ability to successfully complete a given task later in a ‘single-life’ trial, during which there are no human-provided resets. The robot is provided with a diverse set of prior behaviors trained through episodic RL, and the single life poses a sequence of new situations.
@@
The main contribution of this paper is a simple algorithm for autonomous, deployment-time adaptation to novel scenarios. At a given state, with the additional cross-entropy regularizer, ROAM can constrain each behavior’s value to be lower than the value of behaviors for which that state appears more frequently. Consequently, our method incentivizes each behaviors to differentiate between familiar and unfamiliar states, allowing ROAM to better recognize when a behavior will be useful. We conduct experiments on both simulated locomotion tasks and on a real Go1 quadruped robot. In simulation, our method completes the deployment task more than two times faster on average than existing methods, including two prior methods designed for fast adaptation. We also empirically analyze how the additional cross-entropy term in the loss function of ROAM contributes to more successful utilization of the prior behaviors. Furthermore, ROAM enables the Go1 robot to adapt on-the-go to various OOD situations without human interventions or supervision in the real world. With ROAM, the robot can successfully pull heavy luggage, pull loads with dynamic weights, and even slide forward with two roller skates on its front feet, even though it never encountered loads or wore roller skates during training.
@@ 2 Related Work
We consider the problem of enabling an agent to act robustly when transferring to unstructured test-time conditions that are unknown at train-time. One common instantiation of this problem is transfer to different dynamics, e.g., in order to transfer policies trained in simulation to the real world. A popular approach in achieving transfer under dynamics shift is domain randomization, i.e., randomizing the dynamics during training [cutler2014reinforcement, rajeswaran2016epopt, sadeghi2016cad2rl, tobin2017domain, peng2018sim, tan2018sim, yu2019sim, akkaya2019solving, xie2021dynamics, margolis2022rapid, haarnoja2023learning] to learn a robust policy. Our approach is similar in that it takes advantage of different MDPs during training; however, a key component of ROAM is to leverage and modulate diverse skills rather than a single, robust policy. We find in Section 5 that challenging test-time scenarios may require distinctly different behaviors at different times, and we design our method to be robust to those heterogeneous conditions.
@@
A class of methods that use domain randomization has also utilized different ‘behaviors’ instead of a one-size-fits-all policy. Especially effective in locomotion applications, these methods involve training policies that exhibit different behavior when conditioned on dynamics parameters, then distilling these policies into one that can be deployed in target domains where this information is not directly observable. The train-time supervision can come in the form of the parameter values [yu2017preparing, ji2022concurrent] or a learned representation of them [lee2020learning, kumar2021rma]. Thereafter, there are several ways prior work have explored utilizing online data to identify which behavior is appropriate on-the-fly, e.g., using search in latent space [yu2019sim, peng2020learning, yu2020learning], or direct inference using proprioceptive history [lee2020learning, kumar2021rma, fu2023deep], or prediction based on egocentric depth [miki2022learning, agarwal2023legged, zhuang2023robot, yang2023neural]. In this work, we do not rely on domain-specific information nor external supervision for when particular pre-trained behaviors are useful. Moreover, in contrast to many of the above works, we focus on solving tasks that may be OOD for all prior behaviors individually. By re-evaluating and potentially switching behaviors at every timestep, ROAM can take the most useful parts of different behaviors to solve such tasks, as we find in Section 5.
@@
Meta-RL is another line of work that achieves rapid adaptation without privileged information by optimizing the adaptation procedure during training [wang2016learning, duan2016rl, finn2017model, nagabandi2018learning, houthooft2018evolved, rothfuss2018promp, rusu2018meta, mendonca2020meta, song2020rapidly] to be able to adapt quickly to a new situation at test-time. Meta-RL and the aforementioned domain randomization-based approaches entangle the training processes with the data collection, requiring a lot of online samples that are collected in a particular way for pre-training. The key conceptual difference in our approach is that ROAM is agnostic to how the pre-trained policies and value functions are obtained. Moreover, while meta-RL methods often use hundreds of pre-training tasks, or more, our approach can provide improvements in new situations even with a relatively small set of pre-trained behaviors (e.g. just 4 different behaviors improve performance in Section 5). Other transfer learning approaches adapt the weights of the policy to a new environment or task, either through rapid zero-shot adaptation [hansen2020self, yoneda2021invariance, chen2022you] or through extended episodic online training [khetarpal2020towards, rusu2016progressive, eysenbach2020off, xie2020deep, xie2021lifelong]. Unlike these works, we focus on adaptation within a single episode to a variety of different situations.
@@
Another rich body of work considers how to combine prior behaviors to solve long-horizon tasks, and some of these works also focus on learning or discovering useful skills [gregor2016variational, achiam2018variational, eysenbach2018diversity, nachum2018near, sharma2019dynamics, baumli2021relative, laskin2022cic, park2023predictable]. Many of these methods involve training a high-level policy that learns to compose learned skills into long-horizon behaviors [bacon2017option, peng2019mcp, lee2019learning, sharma2020learning, strudel2020learning, nachum2018data, chitnis2020efficient, pertsch2021guided, dalal2021accelerating, nasiriany2022augmenting]. We show in Section 5 that training such a high-level policy is not needed for effective on-the-go behavior selection. Moreover, our work does not focus on where the behaviors come from. Instead, we focus on quickly selecting and adapting the most suitable skill in an OOD scenario, without requiring an additional online training phase to learn a hierarchical controller.
@@ 3 Preliminaries
In this section, we describe some preliminaries and formalize our problem statement. We are given a set of $n$ prior behaviors, where each behavior $i$ is trained through episodic RL for a particular MDP $\mathcal{M}_{i}=(\mathcal{S},\mathcal{A},\mathcal{P}_{i},\mathcal{R}_{i},\rho_{0},\gamma)$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the agent’s action space, $\mathcal{P}_{i}(s_{t+1}|s_{t},a_{t})$ represents the environment’s transition dynamics, $\mathcal{R}_{i}:\mathcal{S}\rightarrow\mathbb{R}$ indicates the reward function, $\rho_{0}:\mathcal{S}\rightarrow\mathbb{R}$ denotes the initial state distribution, and $\gamma\in[0,1)$ denotes the discount factor. Each of the $n$ MDPs $\mathcal{M}_{i}$ has potentially different dynamics and reward functions $\mathcal{P}_{i}$ and $\mathcal{R}_{i}$, often leading to different state visitation distributions. Each behavior corresponds to a policy $\pi_{i}$ and a value function $V_{i}$ as well as a buffer of trajectories $\tau\in\mathcal{D}_{i}$ collected during this prior training and relabeled with the reward $\mathcal{R}_{\text{target}}$ from the target MDP. At test time, the agent interacts with a target MDP defined by $\mathcal{M}_{\text{target}}=(\mathcal{S},\mathcal{A},\mathcal{P}_{\text{target}},\mathcal{R}_{\text{target}},\rho_{0},\gamma)$, which presents an aspect of novelty not present in any of the prior MDPs, in the form of new dynamics $\mathcal{P}_{\text{target}}(s_{t+1}\mid s_{t},a_{t})$, which may change over the course of the test-time trial. We operate in a single-life deployment setting [chen2022you] that aims to maximize $J=\sum_{t=0}^{h}\gamma^{t}\mathcal{R}(s_{t})$, where $h$ is the trial horizon, which may be $\infty$. The agent needs to complete the desired task in this target MDP in a single life without any additional supervision or human intervention by effectively selecting and adapting the prior behaviors to the situation at hand.
@@
In actor-critic frameworks, a separate policy is trained to maximize Q-values. These algorithms alternate between policy evaluation, which involves the Bellman operator $B^{\pi}Q=r+\gamma P^{\pi}Q$, and policy improvement, where the policy $\pi(a|s)$ is updated to maximize expected Q-values.
We use a state-of-the-art off-policy actor-critic algorithm RLPD [ball2023rlpd] as our base algorithm for pre-training and fine-tuning, as it has been shown to be effective especially in the latter regime. RLPD builds on regularized soft actor-critic (SAC) [haarnoja2018soft], using Layer Normalization [ba2016layer] in particular.
@@ 4 Robust Autonomous Modulation
We now present our method, Robust Autonomous Modulation (ROAM), which fine-tunes value functions with an additional loss and provides a mechanism for choosing among them, so that at deployment time, the agent can quickly react to its current situation at every timestep by honing in on the most appropriate behavior from its prior behaviors. Our key observation is that with proper regularization, value functions provide a good indication of how well different behaviors will perform, so we can leverage them to quickly identify appropriate behaviors in a given situation, which circumvents the need to learn a separate meta-controller or adaptation module.
@@
The core idea of our method is to directly utilize the expressive power of value functions for adaptive behavior selection, as they inherently contain detailed information about potential rewards associated with different states for each behavior. We propose to use value functions of the behaviors to select the most appropriate behavior for a given state—namely, at each timestep during deployment, choosing one of the behaviors with the high values at that state. Since we already have access to value functions from pre-training the behaviors, this approach does not require any additional training or data collection. Using value functions as the proxy selector also gives much more versatility to the selection mechanism, which can be flexibly controlled on the go by updating the value functions of different behaviors. However, naively using the pre-trained value functions may not lead to high-reward behaviors, due to overestimation of the value functions on new states, as the pre-trained Q-functions may not generalize well to OOD situations. Recent studies in offline RL, for example, have observed an overestimation bias [levine2020offline] due to OOD training-time actions, which can lead to poor performance when deploying the policy in a new environment. To mitigate these issues, works in offline RL have proposed a number of various modifications aimed at regularizing either the policy or the value function [kumar2020conservative, yu2020mopo, yu2021combo]. Although our setting is different, as we deal with OOD states, we face a similar problem of poor generalization of the value functions of the prior behaviors to unfamiliar situations. In the following section, we describe how we can conservatively regularize the value functions to improve their generalization.
@@
Given a set of prior behaviors $\mathcal{B}$ with policies $\pi_{i}$ and critics $Q_{i}$, we first fine-tune the value functions of the behaviors with an additional cross-entropy loss on top of the Bellman error that takes in the values at a given state as the logits. Values at a given state $s$ for behavior $i$ are obtained by averaging $Q_{i}(s,a)$ over $N=5$ sampled actions $a\sim\pi_{i}(\cdot\mid s)$. More formally, with each prior data buffer $\mathcal{D}_{i}$, we fine-tune the critic $Q_{i}(s,a)$ of each behavior $i$ with the following update:
@@
where 0 < $\beta$ < 1 is a hyperparameter, $V_{i}(s)=\mathbb{E}_{a\sim\pi_{i}(a|s)}[Q_{i}(s,a)]$ is the average value of behavior $i$ at state $s$, and $\mathcal{D}_{i}$ is a buffer collected by behavior $i$. Consider the derivative of the cross-entropy term with respect to the value functions $V_{i}(s)$ and $V_{k}(s)$, where $i\neq k$, for a state $s$ in $\mathcal{D}_{i}$:
@@
To summarize the full procedure of ROAM, we are given a set of policies $\pi_{i}$ and critics $Q_{i}$, and a set of prior data buffers $\mathcal{D}_{i}$ for each behavior. Alternatively, this can be relaxed and we can assume that we are given a set of prior data buffers $\mathcal{D}_{i}$ for each behavior, and we can train the policies $\pi_{i}$ and critics $Q_{i}$ using these buffers with offline RL. We then fine-tune the value functions of the behaviors with the additional cross-entropy loss term in Equation 1 to obtain the final value functions $V_{i}$.
@@
Then during each timestep at test time, we sample a behavior from the softmax distribution given by the behaviors’ values $V_{i}$ at the current state. Formally, given the current state $s$, we sample an action $a_{i}\sim\pi_{i}(a|s)$ from behavior $i$ with probability proportional to $\exp(V_{i}(s))$. The transition $(s_{t},a_{t},r_{t},s_{t+1})$ is then added to the online buffer $\mathcal{D}^{i}_{\text{online}}$ for behavior $i$ and the critic $V_{i}$ and policy $\pi_{i}$ are fine-tuned using data from $\mathcal{D}^{i}_{\text{online}}$. In this manner, we can choose and adapt the most suitable behavior for a given state on-the-fly. The ROAM fine-tuning objective and single-life deployment are depicted in Figure 2 and the full algorithm is summarized in Algorithms 1 and 2 in Appendix A.2. In Appendix A.1, we provide theoretical analysis of ROAM to show that the additional cross-entropy loss in ROAM will lead to more suitable behaviors being chosen.

### v3

@@ Abstract
To succeed in the real world, robots must cope with situations that differ from those seen during training. We study the problem of adapting on-the-fly to such novel scenarios during deployment, by drawing upon a diverse repertoire of previously-learned behaviors. Our approach, RObust Autonomous Modulation (ROAM), introduces a mechanism based on the perceived value of pre-trained behaviors to select and adapt pre-trained behaviors to the situation at hand. Crucially, this adaptation process all happens within a single episode at test time, without any human supervision. We provide theoretical analysis of our selection mechanism and demonstrate that ROAM enables a robot to adapt rapidly to changes in dynamics both in simulation and on a real Go1 quadruped, even successfully moving forward with roller skates on its feet. Our approach adapts over 2x as efficiently compared to existing methods when facing a variety of out-of-distribution situations during deployment by effectively choosing and adapting relevant behaviors on-the-fly.
Figure 1: On-The-Go Adaptation via Robust Autonomous Modulation (ROAM). An agent will inevitably encounter a wide variety of situations during deployment, and handling such situations may require a variety of different behaviors. We propose Robust Autonomous Modulation (ROAM), which dynamically employs relevant behaviors as different situations arise during deployment.
@@ 1 Introduction
Existing adaptation methods often operate within the standard, episodic training paradigm where the agent is assumed to be reset and have another chance to attempt the task each time (Cully et al., 2015; Song et al., 2020; Julian et al., 2020). However, these are idealized conditions that in practice often rely on human intervention. In the above search-and-rescue example, the robot’s state cannot be arbitrarily restored; during deployment, the robot cannot be repaired, and it is not feasible for a human to fetch it if it gets stuck in a situation it is not equipped to handle. This situation necessitates adapting at test time both quickly and autonomously, to succeed at the task within a single episode. Therefore, we frame our problem setting as an instantiation of single-life deployment (Chen et al., 2022), where the agent is given prior knowledge but evaluated on its ability to successfully complete a given task later in a ‘single-life’ trial, during which there are no human-provided resets. The robot is provided with a diverse set of prior behaviors trained through episodic RL, and the single life poses a sequence of new situations.
@@
The main contribution of this paper is a simple algorithm for autonomous, deployment-time adaptation to novel scenarios. At a given state, with the additional cross-entropy regularizer, ROAM can constrain each behavior’s value to be lower than the value of behaviors for which that state appears more frequently. Consequently, our method incentivizes each behavior to differentiate between familiar and unfamiliar states, allowing ROAM to better recognize when a behavior will be useful. We conduct experiments on both simulated locomotion tasks and on a real Go1 quadruped robot. In simulation, our method completes the deployment task more than two times faster on average than existing methods, including two prior methods designed for fast adaptation. We also empirically analyze how the additional cross-entropy term in the loss function of ROAM contributes to more successful utilization of the prior behaviors. Furthermore, ROAM enables a Go1 robot to adapt on-the-go to various OOD situations without human interventions or supervision in the real world. With ROAM, the robot can successfully pull heavy luggage, pull loads with dynamic weights, and even slide forward with two roller skates on its front feet, even though it never encountered loads or wore roller skates during training.
@@ 2 Related Work
We consider the problem of enabling an agent to act robustly when transferring to unstructured test-time conditions that are unknown at train-time. One common instantiation of this problem is transfer to different dynamics, e.g., in order to transfer policies trained in simulation to the real world. A popular approach in achieving transfer under dynamics shift is domain randomization, i.e., randomizing the dynamics during training (Cutler et al., 2014; Rajeswaran et al., 2016; Sadeghi & Levine, 2016; Tobin et al., 2017; Peng et al., 2018; Tan et al., 2018; Yu et al., 2019; Akkaya et al., 2019; Xie et al., 2021; Margolis et al., 2022; Haarnoja et al., 2023) to learn a robust policy. Our approach is similar in that it takes advantage of different MDPs during training; however, a key component of ROAM is to leverage and modulate diverse skills rather than a single, robust policy. We find in Section 5 that challenging test-time scenarios may require distinct behaviors at different times, and we design our method to be robust to those heterogeneous conditions.
@@
Another class of methods involve training policies that exhibit different behavior when conditioned on dynamics parameters, then distilling these policies into one that can be deployed in target domains where this information is not directly observable. The train-time supervision can come in the form of the parameter values (Yu et al., 2017; Ji et al., 2022) or a learned representation of them (Lee et al., 2020; Kumar et al., 2021). Thereafter, there are several ways prior work have explored utilizing online data to identify which behavior is appropriate on-the-fly, e.g., using search in latent space (Yu et al., 2019; Peng et al., 2020; Yu et al., 2020b), or direct inference using proprioceptive history (Lee et al., 2020; Kumar et al., 2021; Fu et al., 2022), or prediction based on egocentric depth (Miki et al., 2022; Agarwal et al., 2022; Zhuang et al., 2023; Yang et al., 2023). In this work, we do not rely on domain-specific information nor external supervision for when particular pre-trained behaviors are useful. Moreover, in contrast to many of the above works, we focus on solving tasks that may be OOD for all prior behaviors individually.
@@
Meta-RL is another line of work that achieves rapid adaptation without privileged information by optimizing the adaptation procedure during training (Wang et al., 2016; Duan et al., 2016; Finn et al., 2017; Nagabandi et al., 2018; Houthooft et al., 2018; Rothfuss et al., 2018; Rusu et al., 2018; Mendonca et al., 2020; Song et al., 2020) to be able to adapt quickly to a new situation at test-time. Meta-RL and the aforementioned domain randomization-based approaches entangle the training processes with the data collection, requiring a lot of online samples that are collected in a particular way for pre-training. The key conceptual difference in our approach is that ROAM is agnostic to how the pre-trained policies and value functions are obtained. Moreover, while meta-RL methods often use hundreds of pre-training tasks, or more, our approach can provide improvements in new situations even with a relatively small set of pre-trained behaviors (e.g. just 4 different behaviors improve performance in Section 5). Other transfer learning approaches adapt the weights of the policy to a new environment or task, either through rapid zero-shot adaptation (Hansen et al., 2020; Yoneda et al., 2021; Chen et al., 2022) or through extended episodic online training (Khetarpal et al., 2020; Rusu et al., 2016; Eysenbach et al., 2020; Xie et al., 2020; Xie & Finn, 2021). Unlike these works, we focus on adaptation within a single episode to a variety of different situations.
@@
Another rich body of work considers how to combine prior behaviors to solve long-horizon tasks, and some of these works also focus on learning or discovering useful skills (Gregor et al., 2016; Achiam et al., 2018; Eysenbach et al., 2018; Nachum et al., 2018a; Sharma et al., 2019; Baumli et al., 2021; Laskin et al., 2022; Park & Levine, 2023). Many of these methods involve training a high-level policy that learns to compose learned skills into long-horizon behaviors (Bacon et al., 2017; Peng et al., 2019; Lee et al., 2019; Sharma et al., 2020; Strudel et al., 2020; Nachum et al., 2018b; Chitnis et al., 2020; Pertsch et al., 2021; Dalal et al., 2021; Nasiriany et al., 2022). We show in Section 5 that such a high-level policy may not be effective for on-the-go behavior selection. Moreover, our work does not focus on where the behaviors come from – they could be produced by these prior methods, or their rewards could be specified manually. Instead, we focus on quickly selecting and adapting the most suitable skill in an OOD scenario, without requiring an additional online training phase to learn a hierarchical controller.
@@ 3 Preliminaries
In this section, we describe some preliminaries and formalize our problem statement. We are given a set of $n$ prior behaviors, where each behavior $b_{i}$ is trained through episodic RL for a particular MDP $\mathcal{M}_{i}=(\mathcal{S},\mathcal{A},\mathcal{P}_{i},\mathcal{R}_{i},\rho_{0},\gamma)$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the agent’s action space, $\mathcal{P}_{i}(s_{t+1}|s_{t},a_{t})$ represents the environment’s transition dynamics, $\mathcal{R}_{i}:\mathcal{S}\rightarrow\mathbb{R}$ indicates the reward function, $\rho_{0}:\mathcal{S}\rightarrow\mathbb{R}$ denotes the initial state distribution, and $\gamma\in[0,1)$ denotes the discount factor. Each of the $n$ MDPs $\mathcal{M}_{i}$ has potentially different dynamics and reward functions $\mathcal{P}_{i}$ and $\mathcal{R}_{i}$, often leading to different state visitation distributions. Each behavior corresponds to a policy $\pi_{i}$ and a value function $V_{i}$ as well as a buffer of trajectories $\tau\in\mathcal{D}_{i}$ collected during this prior training and relabeled with the reward $\mathcal{R}_{\text{target}}$ from the target MDP. At test time, the agent interacts with a target MDP defined by $\mathcal{M}_{\text{target}}=(\mathcal{S},\mathcal{A},\mathcal{P}_{\text{target}},\mathcal{R}_{\text{target}},\rho_{0},\gamma)$, which presents an aspect of novelty not present in any of the prior MDPs, in the form of new dynamics $\mathcal{P}_{\text{target}}(s_{t+1}\mid s_{t},a_{t})$, which may change over the course of the test-time trial. We operate in a single-life deployment setting (Chen et al., 2022) that aims to maximize $J=\sum_{t=0}^{h}\gamma^{t}\mathcal{R}(s_{t})$, where $h$ is the trial horizon, which may be $\infty$. The agent needs to complete the desired task in this target MDP in a single life without any additional supervision or human intervention by effectively selecting and adapting the prior behaviors to the situation at hand.
@@
In actor-critic frameworks, a separate policy is trained to maximize Q-values. These algorithms alternate between policy evaluation, which involves the Bellman operator $B^{\pi}Q=r+\gamma P^{\pi}Q$, and policy improvement, where the policy $\pi(a|s)$ is updated to maximize expected Q-values. We use a state-of-the-art off-policy actor-critic algorithm RLPD (Ball et al., 2023) as our base algorithm for pre-training and fine-tuning, which builds on regularized soft actor-critic (Haarnoja et al., 2018).
@@ 4 Robust Autonomous Modulation
We now present our method, Robust Autonomous Modulation (ROAM), which fine-tunes value functions with an additional loss and provides a mechanism for choosing among them, so that at deployment time, the agent can quickly react to its current situation at every timestep by honing in on the most appropriate behavior from its prior behaviors. Our key observation is that with proper regularization, value functions provide a good indication of how well different behaviors will perform, so we can leverage them to quickly identify appropriate behaviors, which circumvents the need to learn a separate meta-controller or adaptation module.
@@
The core idea of our method is to directly utilize the expressive power of value functions for adaptive behavior selection, as they inherently contain detailed information about potential rewards associated with different states for each behavior. We propose to use value functions of the behaviors to select the most appropriate behavior for a given state—namely, at each timestep during deployment, choosing one of the behaviors with the high values at that state. Since we already have access to value functions from pre-training the behaviors, this approach does not require any additional training or data collection. Using value functions as the proxy selector also gives much more versatility to the selection mechanism, which can be flexibly controlled on the go by updating the value functions of different behaviors. However, naively using the pre-trained value functions may not lead to high-reward behaviors, due to overestimation of the value functions on new states, as the pre-trained Q-functions may not generalize well to OOD situations. Recent studies in offline RL, for example, have observed an overestimation bias (Levine et al., 2020) due to OOD training-time actions, which can lead to poor performance when deploying the policy in a new environment. To mitigate these issues, works in offline RL have proposed a number of various modifications aimed at regularizing either the policy or the value function (Kumar et al., 2020; Yu et al., 2020a; 2021). Although our setting is different, as we deal with OOD states, we face a similar problem of poor generalization of the value functions of the prior behaviors to unfamiliar situations. In the following section, we describe how we can conservatively regularize the value functions to improve their generalization.
@@
Given a set of prior behaviors $\mathcal{B}$ with policies $\pi_{i}$ and critics $Q_{i}$, we first fine-tune the value functions of the behaviors with an additional cross-entropy loss on top of the Bellman error that takes in the values at a given state as the logits. Values at a given state $s$ for behavior $b_{i}$ are obtained by averaging $Q_{i}(s,a)$ over $N=5$ sampled actions $a\sim\pi_{i}(\cdot\mid s)$. More formally, with each prior data buffer $\mathcal{D}_{i}$, we fine-tune the critic $Q_{i}(s,a)$ of each behavior $b_{i}$ with the following update:
@@
where $10<\beta<$ 1 is a hyperparameter, $V_{i}(s)=\mathbb{E}_{a\sim\pi_{i}(a|s)}[Q_{i}(s,a)]$ is the average value of behavior $b_{i}$ at state $s$, and $\mathcal{D}_{i}$ is a replay buffer collected by behavior $b_{i}$. Consider the derivative of the cross-entropy term with respect to the value functions $V_{i}(s)$ and $V_{k}(s)$, where $i\neq k$, for a state $s$ in $\mathcal{D}_{i}$:
@@
To summarize the full procedure of ROAM, we are given a set of policies $\pi_{i}$ and critics $Q_{i}$, and a set of prior data buffers $\mathcal{D}_{i}$ for each behavior, which have been relabeled with the target MDP reward. Alternatively, this can be relaxed and we can assume that we are given a set of prior data buffers $\mathcal{D}_{i}$ for each behavior, and we can train the policies $\pi_{i}$ and critics $Q_{i}$ using these buffers with offline RL.
We then fine-tune the value functions of the behaviors, $V_{i,\text{orig}}$, by incorporating the additional cross-entropy loss term from Equation 1, yielding updated value functions $V_{i}$ for behavior selection. We retain copies of the original value functions to ensure that, if we fine-tune the policies at test time, they remain unaffected by the newly learned value functions. This prevents the ROAM cross-entropy loss from influencing policy optimization, ensuring that the policy updates rely only on their original training objectives.
At each timestep during test time, a behavior is sampled from a softmax distribution over the behaviors’ values $V_{i}$ at the current state. Specifically, given the current state $s$, an action $a_{i}\sim\pi_{i}(a|s)$ is drawn from behavior $b_{i}$, with selection probability proportional to $\exp(V_{i}(s))$. The transition $(s_{t},a_{t},r_{t},s_{t+1})$ is then stored in the online buffer $\mathcal{D}^{i}_{\text{online}}$ for behavior $b_{i}$, and both the critic $V_{i}$, policy $\pi_{i}$, and the original value function $V_{i,\text{orig}}$ are fine-tuned using data from $\mathcal{D}^{i}_{\text{online}}$. Importantly, the fine-tuned value functions influence only the selection of behaviors and do not modify the policies themselves, ensuring that policy optimization remains unbiased during test-time adaptation.
@@
The ROAM fine-tuning objective and single-life deployment are depicted in Figure 2 and the full algorithm is summarized in Algorithms 1 and 2 in Appendix A.2. In the next subsection, we additionally provide theoretical analysis of ROAM to show that the additional cross-entropy loss in ROAM will lead to more suitable behaviors being chosen.
4.1 Theoretical Analysis
Next, we theoretically analyze ROAM to show that the additional cross-entropy loss in ROAM will lead to more suitable behaviors being chosen at each timestep. In particular, ROAM rescales the value functions of the behaviors so that they are less likely to overestimate in states that are out of distribution for that behavior. Our main result, given in Theorem 4.1, is that with ROAM, for some weight $\beta>0$ on the cross-entropy term, at a given state, ROAM constrains each behavior’s value to be lower than the value of behaviors for which that state appears more frequently. This theorem gives us conservative generalization by reducing value overestimation in unfamiliar states–specifically, our chosen behavior will not have worse performance than the most familiar behavior.
This theoretical result holds for all local extrema. Under perfect optimization, ROAM will reach a local minimum at convergence and the result would hold in this case. Our analysis implicitly assumes a simplified tabular setting in the absence of approximation and sampling errors, as done in prior theoretical analysis of deep RL algorithms, e.g. some of the analysis in (Kumar et al., 2020). We take the full gradient of the TD error in the proof although in practice, only the semi-gradient is used during optimization. Regardless, the optimal value function is still a minimum of $\mathcal{L}_{\text{fine-tune}}$ and should satisfy the Bellman equation in the loss function at convergence. We also assume that for all states $s$, we have $V_{\text{freq}}(s)-V_{i}(s)=\left(R_{\text{freq}}(s)+\gamma E_{s^{\prime}}V_{\text{freq}}(s^{\prime})\right)-\left(R_{i}(s)+\gamma E_{s^{\prime}}V_{i}(s^{\prime})\right)<\eta<\infty$. A full proof of the statement in this section is presented in Appendix A.1.
Theorem 4.1.
Let $p_{i}(s)$ denote the state visitation probability for a behavior $b_{i}$ at state $s$. For any state $s$ that is out of distribution for behavior $b_{i}$ and is in distribution for another behavior $b_{j}$, i.e. $p_{i}(s)\ll p_{j}(s)$, if $01>\beta>$ 0 is chosen to be large enough, then the value of behavior $b_{i}$ learned by ROAM will be bounded above compared to value of behavior $b_{j}$, i.e. $V_{i}(s)\leq V_{j}(s)$.
As a result, for any states $s$ that are out of distribution for behavior $b_{i}$ and in distribution for a different behavior $b_{j}$, if we choose $\beta$ large enough, the value learned will not overestimate the value compared to behavior $b_{j}$. Thus, at each time step, if one or more behaviors are familiar with the current state, the performance of the chosen behavior will not be much worse than its value function estimate. In this manner, ROAM adjusts value estimates based on degree of familiarity, mitigating overestimation risks. The ability to adjust the $\beta$ parameter offers a flexible framework to optimize for the behavior with the highest value at a given state, which will be at least as suitable as the most familiar behavior.
In the next section, we find empirically that after fine-tuning with the additional cross-entropy loss, ROAM is able to effectively select a relevant behavior for a given state on-the-fly, leading to robust and fast adaptation to OOD situations.